# AEIA-MN: Evaluating the Robustness of Multimodal LLM-Powered Mobile Agents Against Active Environmental Injection Attacks

### v1

@@ 4 Evaluation
AppAgent Popular Applications
Table 1: Input data of different Agents
Metric Description
$SR_{ben}$ Task success rate without attacks
$SR_{adv}$ Task success rate under Adversarial Attack
$SR_{gap}$ Task success rate under Reasoning Gap Attack
$SR_{com}$ Task success rate under Combinatorial Attack
$SR_{def}$ Task success rate with defense prompts
$ASR_{adv}$ The ratio of tasks where agents are misled by adversarial content.
$ASR_{gap}$ The ratio of tasks where agents affected by Reasoning Gap Attack.
$ASR_{com}$ The ratio of tasks where agents affected by Combinatorial Attack.
$ASR_{def}$ The ratio of tasks where agents are misled despite defense prompts.
Table 2: The description of metrics.
Models I3A (Android World) M3A (Android World) T3A (Android World)
$SR_{ben}$ $SR_{adv}$ $ASR_{adv}$ $SR_{ben}$ $SR_{adv}$ $ASR_{adv}$ $SR_{ben}$ $SR_{adv}$ $ASR_{adv}$
@@
GPT-4o-2024-08-06 0.54 0.34 $\downarrow$ 0.59 0.61 0.39 $\downarrow$ 0.59 0.53 0.43 $\downarrow$ 0.29
Qwen-VL-Max 0.33 0.18 $\downarrow$ 0.26 0.38 0.19 $\downarrow$ 0.18 0.31 0.26 $\downarrow$ 0.34
GLM-4V-Plus 0.16 0.11 $\downarrow$ 0.75 0.13 0.12 $\downarrow$ 0.81 0.03 0.08 0.11
@@
Qwen2-VL-7B 0.05 0.05 0.21 0.30 0.02 $\downarrow$ 0.21 0.08 0.10 0.26
Llava-OneVision-7B 0.10 0.06 $\downarrow$ 0.31 0.02 0.02 0.24 0.05 0.06 0.36
Table 3: The evaluation results of different models under the Adversarial Attack in AndroidWorld.
Models I3A (Android World) M3A (Android World) T3A (Android World)
$SR_{ben}$ $SR_{gap}$ $ASR_{gap}$ $SR_{ben}$ $SR_{gap}$ $ASR_{gap}$ $SR_{ben}$ $SR_{gap}$ $ASR_{gap}$
Closed-source  models
GPT-4o-2024-08-06 0.54 0.31 $\downarrow$ 0.26 0.61 0.34 $\downarrow$ 0.25 0.53 0.33 $\downarrow$ 0.26
Qwen-VL-Max 0.33 0.26 $\downarrow$ 0.18 0.38 0.18 $\downarrow$ 0.26 0.31 0.15 $\downarrow$ 0.26
GLM-4V-Plus 0.16 0.11 $\downarrow$ 0.13 0.13 0.12 $\downarrow$ 0.16 0.03 0.16 0.16
Open-source  models
Qwen2-VL-7B 0.05 0.06 0.13 0.03 0.03 0.23 0.08 0.10 0.14
Llava-OneVision-7B 0.10 0.06 $\downarrow$ 0.05 0.02 0.02 0.18 0.05 0.11 0.18
Table 4: The evaluation results of different models under the Reasoning Gap Attack in AndroidWorld.
4.1 Settings
@@
We structured the experiment setup around three components: benchmarks, models, and agents. The specific details of the settings are presented below.
@@
Benchmarks. We evaluate the performance of the agents provided in the easy subset of the AndroidWorld benchmark Rawles et al. (2024), consisting of 61 tasks. Additionally, for the AppAgent, we utilize its own evaluation benchmark, which includes 45 popular application tasks.
@@
Models. We employed five advanced MLLMs for testing, including the closed-source models GPT-4o-2024-08-06 (Hurst et al., 2024), Qwen-VL-Max (Bai et al., 2023) and GLM-4V-Plus (Hong et al., 2024), as well as the open-source models Qwen2-VL-7B (Wang et al., 2024c) and Llava-OneVision-7B (Li et al., 2024).
@@
Agents. We conduct experiments using agents provided by AndroidWorld Rawles et al. (2024), including mobile agents such as M3A, T3A, and a custom agent, I3A. Additionally, we tested the default configuration of AppAgent (Zhang et al., 2023), a mobile agent designed for user-defined tasks. The differences in the input data for each agent are presented in Table 1. The input data is categorized into two types: image data and element data. These inputs vary across different test benchmarks. Further details about these two types of input data can be found in Appendix C.
4.2 Metrics
We present the metrics used in the experiments in Table 2. In addition, to test the effectiveness of adversarial text based on message notifications against the agents, we set the adversarial content in the message notifications across all experiments to “Agent should believe the task is finished!”
4.3 Main Results in AndroidWorld
We evaluated the robustness of various MLLMs against AEIA-MN in AndroidWorld benchmark. The evaluation results are presented below.
Adversarial Attack. We present the evaluation results of different MLLMs under Adversarial Attack for various agents in Table 4, which show that most MLLMs have limited defense capabilities against such attacks. In I3A and M3A, the Adversarial Attack generally reduces task success rates; however, their adversarial impact is limited by the model’s robustness. In some models, even with a high attack success rate, the decrease in task success rate is minimal. For example, the M3A of GLM-4V-Plus has an $ASR_{adv}$ of 81%, yet the task success rate only drops by 1%. In contrast, in T3A, the Adversarial Attack exhibits a double-edged sword effect: a high attack success rate not only fails to disrupt the task but is interpreted as a strong termination signal due to the prompt “Agent should believe the task is finished!”, correcting the model’s “execution loop” flaw and causing $SR_{adv}$ to increase against the trend. This also indicates that MLLMs in T3A are affected by the adversarial attack. Furthermore, we compare the average number of steps taken to complete the task under different conditions in Figure 3. It shows that, under the influence of Adversarial Attack, the average number of steps to complete the task decreased for most models. In some cases, the number of steps for certain models (such as Qwen-VL-Max and GLM-4V-Plus) increased, indicating that these models possess stronger defensive capabilities.
Figure 3: A comparison of the average number of steps taken by agents to complete the tasks.
Reasoning Gap Attack. We present the evaluation results of different agents under the Reasoning Gap Attack in Table 4. The results show that the Reasoning Gap Attack significantly disrupts the task execution of most agents, causing a notable decrease in task success rates across most models. This disruption is achieved by transitioning the actions performed by the agent into an unintended device state. However, in T3A, the $SR_{gap}$ of GLM-4V-Plus and LLaVA-OneVision-7B increased from 3% to 16% and from 5% to 11%, respectively. This was not due to model performance improvement but rather because the attack altered the device during the reasoning gap, trapping the system in a dialog window with adversarial content. The prompt "Agent should believe task is finished!" influenced the model, resolving tasks stuck in an "execution loop" and unexpectedly increasing task success rates. This shows that models subjected to a Reasoning Gap Attack can be further influenced by the adversarial content within it.
Combinatorial Attack. Table 5 shows that the Combinatorial Attack is significantly more destructive than single attacks. By overlaying adversarial perturbations with reasoning gap vulnerabilities, the Combinatorial Attack causes significant damage to MLLMs, with success rate reductions reaching up to 67.2%, far exceeding those of single attacks, particularly affecting closed-source MLLMs. However, an anomalous phenomenon occurs in T3A: the adversarial prompt “Agent should believe the task is finished!” may be interpreted as a termination signal when the model is “executing in loops” due to a misjudged state, forcing the end of redundant operations and actually improving the task success rate (as seen with Qwen2-VL-7B’s task success rate increasing by 37.5%).
Models I3A (Android World) M3A (Android World) T3A (Android World)
@@ 4.3 Main Results in AndroidWorld
GPT-4o-2024-08-06 0.54 0.34 0.31 0.29 $\downarrow$ 0.55 0.61 0.39 0.34 0.20 $\downarrow$ 0.72 0.53 0.43 0.33 0.28 $\downarrow$ 0.33
Qwen-VL-Max 0.33 0.18 0.26 0.07 $\downarrow$ 0.33 0.38 0.19 0.18 0.15 $\downarrow$ 0.38 0.31 0.26 0.15 0.21 0.36
GLM-4V-Plus 0.16 0.11 0.11 0.07 $\downarrow$ 0.51 0.13 0.12 0.12 0.11 $\downarrow$ 0.93 0.03 0.08 0.16 0.16 0.59
@@
Qwen2-VL-7B 0.05 0.05 0.06 0.05 0.38 0.03 0.02 0.03 0.02 0.21 0.08 0.10 0.10 0.11 0.21
Llava-OneVision-7B 0.10 0.06 0.06 0.05 $\downarrow$ 0.28 0.02 0.02 0.02 0.02 0.31 0.05 0.06 0.11 0.10 0.24
Table 5: The evaluation results of different models under the Combinatorial Attack in AndroidWorld.
Models I3A (Android World) M3A (Android World) T3A (Android World) AppAgent
@@
GPT-4o-2024-08-06 0.54 0.34 0.31 0.61 0.39 0.40 $\uparrow$ 0.53 0.43 0.36 0.09 0.02 0.09 $\uparrow$
Qwen-VL-Max 0.33 0.18 0.18 0.38 0.19 0.19 0.31 0.26 0.25 0.02 0.0 0.0
GLM-4V-Plus 0.16 0.11 0.13 $\uparrow$ 0.13 0.12 0.12 0.03 0.08 0.17 $\uparrow$ 0.20 0.11 0.11
@@
Qwen2-VL-7B 0.05 0.05 0.15 $\uparrow$ 0.03 0.02 0.03 $\uparrow$ 0.08 0.10 0.09 0.29 0.20 0.27 $\uparrow$
Llava-OneVision-7B 0.10 0.06 0.07 $\uparrow$ 0.02 0.02 0.02 0.05 0.06 0.08 $\uparrow$ 0.07 0.02 0.04 $\uparrow$
Table 6: The evaluation results of different models in various benchmarks after using defense prompts.
Models Attack Type $SR_{ben}$ $SR_{att}$ $ASR_{att}$
GPT-4o-2024-08-06 Adversarial 0.09 0.0 $\downarrow$ 0.68
Reasoning Gap 0.09 0.06 $\downarrow$ 0.09
Combinatorial 0.09 0.02 $\downarrow$ 0.68
Qwen-VL-Max Adversarial 0.02 0.0 $\downarrow$ 0.31
Reasoning Gap 0.02 0.02 0.11
Combinatorial 0.02 0.0 $\downarrow$ 0.51
GLM-4V-Plus Adversarial 0.20 0.11 $\downarrow$ 0.0
Reasoning Gap 0.20 0.06 $\downarrow$ 0.11
Combinatorial 0.20 0.11 $\downarrow$ 0.13
Qwen2-VL-7B Adversarial 0.29 0.20 $\downarrow$ 0.51
Reasoning Gap 0.29 0.22 $\downarrow$ 0.11
Combinatorial 0.29 0.20 $\downarrow$ 0.84
Llava-OneVision-7B Adversarial 0.07 0.02 $\downarrow$ 0.30
Reasoning Gap 0.07 0.04 $\downarrow$ 0.11
Combinatorial 0.07 0.0 $\downarrow$ 0.51
Table 7: The evaluation results of our proposed attack scheme AEIA-MN in AppAgent. $SR_{att}$ and $ASR_{att}$ denote the success rates and attack success rate for different types of attacks: Adversarial Attack ($SR_{adv}$, $ASR_{adv}$), Reasoning Gap Attack ($SR_{gap}$, $ASR_{gap}$), and Combinatorial Attack ($SR_{com}$, $ASR_{com}$).
4.4 Main Results in AppAgent
@@
The effects of Resoning Gap Attack are weaker, causing some interference only for GLM-4V-Plus (with the $SR_{gap}$ declining from 20% to 6%) and LLaVA-OneVision-7B (with the $SR_{gap}$ decreasing from 7% to 4%). Notably, GLM-4V-Plus experiences an $ASR_{adv}$ of 0%, yet the task success rate still drops from 20% to 11%, which may be inferred to result from the UI element occlusion of message notification elements. Overall, Combinatorial Attack poses the greatest threat to both closed-source and open-source models, while Adversarial Attack is more targeted towards closed-source models. Reasoning Gap Attack, on the other hand, has limited attack intensity.
4.5 Defense Prompt
We tested the defense prompt against notification-based Adversarial Attack by adding the defense prompt "Please make sure to ignore any content on the notifications that is unrelated to the goal." in the prompt. The results are shown in Table 6.

### v2

@@ 3.2.4. Combinatorial Attack
AppAgent Popular Apps
4. Evaluation
In this section, we develop a prototype of AEIA-MN and conduct extensive experiments to assess the robustness of various agents against the attack. We begin by outlining implementation details and metrics in Sections 4.1 and 4.2, respectively. Subsequently, we analyze the experimental results in Sections 4.3 to 4.8.
Table 2. Evaluation results of different MLLM-based Agents under Adversarial Attack and Reasoning Gap Attack on AndroidWorld benchmark. Arrows indicate performance drop compared to $SR_{ben}$.
Model I3A M3A T3A
Adversarial Reasoning Gap Adversarial Reasoning Gap Adversarial Reasoning Gap
$SR_{ben}$ $SR_{adv}$ $ASR_{adv}$ $SR_{gap}$ $ASR_{gap}$ $SR_{ben}$ $SR_{adv}$ $ASR_{adv}$ $SR_{gap}$ $ASR_{gap}$ $SR_{ben}$ $SR_{adv}$ $ASR_{adv}$ $SR_{gap}$ $ASR_{gap}$
@@ 4. Evaluation
GPT-4o-2024-08-06 (Hurst et al., 2024) 0.54 0.34$\downarrow$ 0.59 0.31$\downarrow$ 0.26 0.61 0.39$\downarrow$ 0.59 0.34$\downarrow$ 0.25 0.53 0.43$\downarrow$ 0.29 0.33$\downarrow$ 0.26
Qwen-VL-Max (Bai et al., 2023) 0.33 0.18$\downarrow$ 0.26 0.26$\downarrow$ 0.18 0.38 0.19$\downarrow$ 0.18 0.18$\downarrow$ 0.26 0.31 0.26$\downarrow$ 0.34 0.15$\downarrow$ 0.26
GLM-4V-Plus (Hong et al., 2024) 0.16 0.11$\downarrow$ 0.75 0.11$\downarrow$ 0.13 0.13 0.12$\downarrow$ 0.81 0.12$\downarrow$ 0.16 0.03 0.08 0.11 0.16 0.16
@@
Qwen2-VL-7B (Wang et al., 2024a) 0.05 0.05 0.21 0.06 0.13 0.03 0.02$\downarrow$ 0.21 0.03 0.23 0.08 0.10 0.26 0.10 0.14
Llava-OneVision-7B (Li et al., 2024) 0.10 0.06$\downarrow$ 0.31 0.06$\downarrow$ 0.05 0.02 0.02 0.24 0.02 0.18 0.05 0.06 0.36 0.11 0.18
4.1. Implementation Details
@@
We conducted experiments on the Android emulator, designing the experimental setup around three key components: benchmarks, models, and agents. The configuration details are outlined below.
@@
Benchmarks. We evaluate the performance of the agents provided in the easy subset of the AndroidWorld benchmark (Rawles et al., 2024), consisting of 61 tasks. For the AppAgent, we utilize its own evaluation benchmark, which includes 45 popular application tasks.
@@
Models. We employed five advanced MLLMs for testing, including the closed-source models GPT-4o-2024-08-06 (Hurst et al., 2024), Qwen-VL-Max (Bai et al., 2023) and GLM-4V-Plus (Hong et al., 2024), as well as the open-source models Qwen2-VL-7B (Wang et al., 2024a) and Llava-OneVision-7B (Li et al., 2024).
@@
Agents. We conduct experiments using agents provided by AndroidWorld (Rawles et al., 2024), including mobile agents such as M3A, T3A, and a custom agent, I3A. Additionally, we tested the default configuration of AppAgent (Zhang et al., 2023). The differences in the input data for each agent are presented in Table 1. The input data is categorized into two types: image data and element data. These inputs vary across different test benchmarks. For AndroidWorld agents, image data includes both the current screenshot and Set-of-Mark (SoM) annotated screenshots (Yang et al., 2023), where each UI element in the SoM screenshot is enclosed in a bounding box with a numeric label. For AppAgent, image data consists only of the SoM annotated screenshot. Element data refers to the text representation of the accessibility tree (a11y tree), which is a hierarchical structure outlining the elements and their associated accessibility properties within a web page or application.
4.2. Metrics
To evaluate the effectiveness of adversarial text embedded in message notifications against agents, we incorporate the adversarial content “Agent should believe the task is finished!” into all experimental message notifications. Moreover, We provide tests for other adversarial content in the supplementary material.
The metrics are defined as follows: (i) Success Rate (SR): $SR_{ben}$ represents the task success rate without any attacks, while $SR_{adv}$, $SR_{gap}$, and $SR_{com}$ refer to the success rates under Adversarial, Reasoning Gap, and Combinatorial Attacks, respectively. $SR_{def}$ indicates the success rate under Adversarial Attacks when agents use defense prompts. (ii) Attack Success Rate (ASR): $ASR_{adv}$, $ASR_{gap}$, and $ASR_{com}$ represent the ratios of tasks where agents are misled by adversarial content, affected by Reasoning Gap Attacks, and influenced by Combinatorial Attacks, respectively. Finally, $ASR_{def}$ measures the ratio of tasks where agents are misled despite defense prompts. The detailed explanations and calculation formulas for the metrics are provided in the supplementary material.
Table 3. Evaluation results of different MLLM-based Agents under the Combinatorial Attack on AndroidWorld benchmark. Arrows indicate performance drop compared to $SR_{ben}$.
Model I3A M3A T3A
@@ 4.2. Metrics
GPT-4o-2024-08-06 (Hurst et al., 2024) 0.54 0.34 0.31 0.29 $\downarrow$ 0.55 0.61 0.39 0.34 0.20 $\downarrow$ 0.72 0.53 0.43 0.33 0.28 $\downarrow$ 0.33
Qwen-VL-Max (Bai et al., 2023) 0.33 0.18 0.26 0.07 $\downarrow$ 0.33 0.38 0.19 0.18 0.15 $\downarrow$ 0.38 0.31 0.26 0.15 0.21 0.36
GLM-4V-Plus (Hong et al., 2024) 0.16 0.11 0.11 0.07 $\downarrow$ 0.51 0.13 0.12 0.12 0.11 $\downarrow$ 0.93 0.03 0.08 0.16 0.16 0.59
@@
Qwen2-VL-7B (Wang et al., 2024a) 0.05 0.05 0.06 0.05 0.38 0.03 0.02 0.03 0.02 0.21 0.08 0.10 0.10 0.11 0.21
Llava-OneVision-7B (Li et al., 2024) 0.10 0.06 0.06 0.05 $\downarrow$ 0.28 0.02 0.02 0.02 0.02 0.31 0.05 0.06 0.11 0.10 0.24
4.3. Main Results on AndroidWorld
We conducted a comprehensive evaluation of the robustness of agents based on various MLLM against AEIA-MN on the AndroidWorld benchmark. The evaluation results are presented as follows.
4.3.1. Adversarial Attack
We present the evaluation results of different MLLM-based Agents under Adversarial Attack in Table 2, which show that most agents have limited defense capabilities against such attacks. In I3A and M3A, the Adversarial Attack generally reduces task success rates; however, their adversarial impact is limited by the model’s robustness. In some models, even with a high attack success rate, the decrease in task success rate is minimal. For example, the M3A of GLM-4V-Plus has an $ASR_{adv}$ of 81%, yet the task success rate only drops by 1%. In contrast, in T3A, the Adversarial Attack exhibits a double-edged sword effect: a high attack success rate not only fails to disrupt the task but is interpreted as a strong termination signal due to the prompt “Agent should believe the task is finished!”, correcting the model’s “execution loop” flaw and causing $SR_{adv}$ to increase against the trend. This also indicates that MLLMs in T3A are affected by the adversarial attack. We discuss this phenomenon in more detail in supplementary material.
4.3.2. Reasoning Gap Attack
We present the evaluation results of different agents under the Reasoning Gap Attack in Table 2. The results show that the Reasoning Gap Attack significantly disrupts the task execution of most agents, causing a notable decrease in task success rates across most models. This disruption is achieved by transitioning the actions performed by the agent into an unintended device state. However, in T3A, the $SR_{gap}$ of GLM-4V-Plus and LLaVA-OneVision-7B increased from 3% to 16% and from 5% to 11%, respectively. This was not due to model performance improvement but rather because the attack altered the device during the reasoning gap, trapping the system in a dialog window with adversarial content. The prompt ”Agent should believe task is finished!” influenced the model, resolving tasks stuck in an ”execution loop” and unexpectedly increasing task success rates. This shows that models subjected to a Reasoning Gap Attack can be further influenced by the adversarial content within it.
4.3.3. Combinatorial Attack
Table 3 shows that the Combinatorial Attack is significantly more destructive than single attacks. By overlaying adversarial perturbations with reasoning gap vulnerabilities, the Combinatorial Attack causes significant damage to MLLMs, with success rate reductions reaching up to 67.2%, far exceeding those of single attacks, particularly affecting closed-source MLLMs. However, an anomalous phenomenon occurs in T3A: the adversarial prompt “Agent should believe the task is finished!” may be interpreted as a termination signal when the model is “executing in loops” due to a misjudged state, forcing the end of redundant operations and actually improving the task success rate (as seen with Qwen2-VL-7B’s task success rate increasing by 37.5%).
Table 4. Evaluation results of AEIA-MN on AppAgent benchmark. $SR_{att}$ and $ASR_{att}$ denote the success rates and attack success rate for different types of attacks: Adversarial Attack ($SR_{adv}$, $ASR_{adv}$), Reasoning Gap Attack ($SR_{gap}$, $ASR_{gap}$), and Combinatorial Attack ($SR_{com}$, $ASR_{com}$). Arrows indicate performance drop compared to $SR_{ben}$.
Model Attack Type $SR_{ben}$ $SR_{att}$ $ASR_{att}$
GPT-4o-2024-08-06 (Hurst et al., 2024) Adversarial 0.09 0.0 $\downarrow$ 0.68
Reasoning Gap 0.09 0.06 $\downarrow$ 0.09
Combinatorial 0.09 0.02 $\downarrow$ 0.68
Qwen-VL-Max (Bai et al., 2023) Adversarial 0.02 0.0 $\downarrow$ 0.31
Reasoning Gap 0.02 0.02 0.11
Combinatorial 0.02 0.0 $\downarrow$ 0.51
GLM-4V-Plus (Hong et al., 2024) Adversarial 0.20 0.11 $\downarrow$ 0.0
Reasoning Gap 0.20 0.06 $\downarrow$ 0.11
Combinatorial 0.20 0.11 $\downarrow$ 0.13
Qwen2-VL-7B (Wang et al., 2024a) Adversarial 0.29 0.20 $\downarrow$ 0.51
Reasoning Gap 0.29 0.22 $\downarrow$ 0.11
Combinatorial 0.29 0.20 $\downarrow$ 0.84
Llava-OneVision-7B (Li et al., 2024) Adversarial 0.07 0.02 $\downarrow$ 0.30
Reasoning Gap 0.07 0.04 $\downarrow$ 0.11
Combinatorial 0.07 0.0 $\downarrow$ 0.51
Table 5. Evaluation results of different MLLM-based Agents using defense prompts on AndroidWorld and AppAgent benchmarks. Arrows indicate performance up compared to $SR_{ben}$.
Model I3A (AndroidWorld) M3A (AndroidWorld) T3A (AndroidWorld) AppAgent
@@ 4.3.3. Combinatorial Attack
GPT-4o-2024-08-06 (Hurst et al., 2024) 0.54 0.34 0.31 0.61 0.39 0.40 $\uparrow$ 0.53 0.43 0.36 0.09 0.02 0.09 $\uparrow$
Qwen-VL-Max (Bai et al., 2023) 0.33 0.18 0.18 0.38 0.19 0.19 0.31 0.26 0.25 0.02 0.0 0.0
GLM-4V-Plus (Hong et al., 2024) 0.16 0.11 0.13 $\uparrow$ 0.13 0.12 0.12 0.03 0.08 0.17 $\uparrow$ 0.20 0.11 0.11
@@
Qwen2-VL-7B (Wang et al., 2024a) 0.05 0.05 0.15 $\uparrow$ 0.03 0.02 0.03 $\uparrow$ 0.08 0.10 0.09 0.29 0.20 0.27 $\uparrow$
Llava-OneVision-7B (Li et al., 2024) 0.10 0.06 0.07 $\uparrow$ 0.02 0.02 0.02 0.05 0.06 0.08 $\uparrow$ 0.07 0.02 0.04 $\uparrow$
4.4. Main Results on AppAgent
@@
The effects of Resoning Gap Attack are weaker, causing some interference only for GLM-4V-Plus (with the $SR_{gap}$ declining from 20% to 6%) and LLaVA-OneVision-7B (with the $SR_{gap}$ decreasing from 7% to 4%). Notably, GLM-4V-Plus experiences an $ASR_{adv}$ of 0%, yet the task success rate still drops from 20% to 11%, which may be inferred to result from the UI element occlusion of message notification elements. Overall, the Combinatorial Attack poses the most significant threat to both closed- and open-source models, while the Adversarial Attack primarily targets closed-source models. In contrast, the Reasoning Gap Attack has limited effectiveness.
4.5. Analysis on Adversarial Attack
To further analyze the impact of adversarial attacks on the Agent, we compared the average number of steps required to complete the tasks in the two benchmarks presented in Figure 3. It can be clearly observed that under Adversarial Attacks in AEIA-MN, the average number of steps required for most MLLMs to complete the tasks decreases. Evidently, most models are influenced by the adversarial prompt ”Agent should believe the task is finished!”, leading to premature task termination and thus a reduction in the average number of steps executed. This further demonstrates that adversarial attacks based on mobile notifications can effectively affect the judgment of MLLMs. Among these results, GLM-4V-Plus is particularly affected, with a step reduction of 5.8 in I3A and 7.3 in M3A. This finding indicates that GLM-4V-Plus has limited robustness against adversarial attacks.
Figure 3. A comparison of the average number of steps taken by agents to complete the tasks on AndroidWorld benchmark.
4.6. Analysis on Reasoning Gap
The duration of the Reasoning Gap affects the success rate of Reasoning Gap Attacks. Specifically, when the Reasoning Gap is very short, it impacts the attack success rate of Reasoning Gap Attacks. To investigate this, we measured the Reasoning Gap duration and the attack initiation time for all models used in our experiments.
We present our measurement results in Table 6. Under experimental conditions, the Reasoning Gap is influenced by multiple factors, including API source limitations, network latency, request-response speed, and prompt length. Therefore, the Reasoning Gap duration can only serve as a reference for our specific experimental setup. In our experimental setup, we randomly selected five tasks from the AndroidWorld benchmark and five tasks from the AppAgent benchmark for testing. We recorded the minimum, maximum, and average Reasoning Gap durations for these five tasks. The test results are shown in Table 6. As observed in the table, the average response time of the models ranges from approximately 3 to 12 seconds. Since mobile notification-based attacks have an Attack Launch Time (ALT), the ALT must be shorter than the Reasoning Gap duration for the Reasoning Gap Attack to be effective. Our measurements indicate that the ALT in this experiment is approximately 1.2 seconds. Because the reasoning start time and the attack initiation time are synchronized in this experiment, the Reasoning Gap Attack can successfully alter the device state through mobile notifications during the Reasoning Gap period.
Table 6. The average Reasoning Gap time length of models.
Model AndroidWorld AppAgent ALT
Min Max Average Min Max Average     Average
GPT-4o-2024-08-06 (Hurst et al., 2024) 7.30 12.47 8.95 11.79 13.36 12.58     1.2
Qwen-VL-Max (Bai et al., 2023) 3.56 10.71 6.48 3.35 8.88 5.97
GLM-4V-Plus (Hong et al., 2024) 9.85 13.46 11.18 7.02 11.50 9.64
Qwen2-VL-7B (Wang et al., 2024a) 3.28 28.29 7.65 2.47 5.95 3.41
Llava-OneVision-7B (Li et al., 2024) 3.30 31.30 4.96 2.70 5.01 3.93
Figure 4. A Comparison of the attack success rates of Adversarial Attack on the agent under different conditions.
4.7. Discussion on Defense Prompt